Forums

We want to share an important update with our Betfair Community users.

Over the coming months, the Betfair Community platform will be winding down. The final active day for the platform will be December 21st. From December 22nd, users will no longer be able to access the platform.

Your discussions, insights, engagement, and, most importantly, feedback have helped shape the platform and guided innovation. Your feedback continues to be important to us, and we will continue to listen via our dedicated feedback channels, as well as through product testing and innovations.

To share your feedback and experience, visit: https://support.betfair.com/app/home

General Betting

Welcome to Live View – Take the tour to learn more
Start Tour
There is currently 1 person viewing this thread.
fatboy1945
20 Dec 09 11:07
Joined:
Date Joined: 18 Dec 02
| Topic/replies: 12 | Blogger: fatboy1945's blog
Can someone tell me how do i get a back price from a lay price.
eg. if i lay a bet at 3.7 what would the equivalent back price be.
I hope this makes sense
Pause • Switch to Standard View Lay odds to back odds
Show More
Loading...
Report Honest Al • December 20, 2009 11:15 AM GMT
fatboy1945,

Divide your original odds into 100, in this case 100/3.7 = 27.027%
The reverse odds would be the other 72.972%
i.e. 100 divided by 72.972 = odds of 1.27
Report [x] These checkboxes suck • December 20, 2009 11:16 AM GMT
pmsl
Report Honest Al • December 20, 2009 11:19 AM GMT
Try again MINUS one point.

100 divided by 2.7 = 37.037%
Reverse odds would be 62.96% = 1.588
Report fatboy1945 • December 20, 2009 11:19 AM GMT
Many thanks honest
Report Contrarian • December 20, 2009 11:25 AM GMT
It's easier just to divide the lay odds, L, by L - 1.

Eg. 3.7 is equivalent to 3.7/2.7 (1.37)
Report Treble_Underscore • December 20, 2009 11:26 AM GMT
1/(3.7-1) = 1/2.7 = 0.3703703

add 1 back on, 1.3703703

reverse of 3.7 certainly isn't 1.27!

3.7 is nearly 11/4 so reverse would be around 4/11 - a rough ready reckoner.
Report scoop6winner. • December 20, 2009 11:30 AM GMT
I was always told never to mix fractions and decimals. but you are correct...
Report Mudflaps • December 20, 2009 11:31 AM GMT
So it's either 1.27, 1.37 or 1.59.

:)
Report JamesBlakesHugeArse • December 20, 2009 11:33 AM GMT
:) unlucky Honest
Post Your Reply
<CTRL+Enter> to submit
Please login to post a reply.

Wonder

Instance ID: 13539
www.betfair.com